A soda manufacturer is interested in determining whether it's bottling machine tends to overfill. Each bottle is supposed to contain 15 ounces of fluid. A random sample of 30 bottles was taken and found that the mean amount of soda of the sample of bottles is 13.4 ounces with a standard deviation of 2.98 ounces. if the manufacturer decides on a significance level od 0.05 test, should the null hypothesis be rejected?


1
Expert's answer
2022-06-16T06:21:22-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le 15"

"H_1:\mu>15"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=29" and the critical value for a right-tailed test is "t_c =1.699127."

The rejection region for this right-tailed test is "R = \{t:t>1.699127\}."

The t-statistic is computed as follows:



"t=\dfrac{11.9-10}{1.8/\sqrt{25}}=\dfrac{13.4-15}{2.98/\sqrt{30}}=-2.9408"


Since it is observed that "t=-2.9408<1.699127=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, "df=29" degrees of freedom, "t=-2.9408" is "p=0.996814," and since "p=0.996814>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is greater than 15, at the "\alpha = 0.05" significance level.